I need help to solve by substitution method for 6m+n=46, m-9n=81. Thanks!
Answer by stanbon(75887)

substitution method for
6m+n=46,
m-9n=81
----------------
m = 9n+81
--------------------
Substitute into the 1st equation and solve for "n":
6(9n+81)+ n = 46
--------
54n + 486 + n = 46
55n = -440
---
n = -8
=================
Substitute into m = 9n+81 to solve for "n":
m = 9(-8)+81
m = 9
===================
Cheers,
Stan H.